# Creating and Benchmarking a Synthetic Dataset for Cloud Optical Thickness Estimation

### v1

@@ 10.2 Fine-tuning with Weaker Labels
Let $\tau^{\text{semi}}$ and $\tau^{\text{opaque}}$ ($0<\tau^{\text{semi}}<\tau^{\text{opaque}}$) denote the semi-transparent and opaque COT thresholds, respectively. Then, we refine a model using a loss $\mathcal{L}$ which satisfies the following criteria. If $p$ denotes the prediction of a pixel which is labeled as
  • ‘clear’, then $p-\tau^{\text{semi}}$ has 0 as regression target for $p>\tau^{\text{semi}}$ and otherwise has no regression target, i.e. $\mathcal{L}(p)=\frac{1}{2}\|\mathrm{ReLU}{}(p-\tau^{\text{semi}})\|^{2}$;
  • ‘opaque cloud’, then $\tau^{\text{opaque}}-p$ has 0 as regression target for $p<\tau^{\text{opaque}}$ and otherwise has no regression target, i.e. $\mathcal{L}(p)=\frac{1}{2}\|\mathrm{ReLU}{}(\tau^{\text{opaque}}-p)\|^{2}$;
  • ‘semi-transparent cloud’, then $\tau^{\text{semi}}-p$ has 0 as regression target for $p<\tau^{\text{semi}}$, $p-\tau^{\text{opaque}}$ has 0 as regression target for $p>\tau^{\text{opaque}}$, and otherwise has no regression target, i.e. $\mathcal{L}(p)=\frac{1}{2}\|\mathrm{ReLU}{}(\tau^{\text{semi}}-p)+\mathrm{ReLU}{}(p-\tau^{\text{opaque}})\|^{2}$.
Table 2: RMSE values on different variants of the test set of our synthetic dataset. Test-xrefers to the test set with $x$added noise. Ensemble methods marginally improve over single-model ones. Models trained with additive input noise yield significantly better average results. Linear regression performs worst by far. Note that for the single-model variants MLP-5, MLP-5-no-noise and Lin-reg, we show the mean RMSE over 10 different network parameter initializations, as well as standard deviations.
@@
Table 3: Results on the test set of KappaZeta. MLP approaches do not perform as well as U-nets, which is expected as U-nets integrate information from other pixels, which MLPs inherently cannot. See also qualitative results in Fig. 2 (right). Among the MLP approaches, model fine-tuning on the KappaZeta training set only slightly improves results (column 4 vs 3), i.e. our synthetic dataset allows for model generalization to real data, while ensembling tuned models negligibly improves results (column 4 vs 2).
@@
12 Experimental Results
13 sec:models
. For COT estimation, we evaluate the following models: (i) MLP-k is a $k$-layer MLP (the main model has $k=5$), where results are averaged among 10 identical $k$-layer MLP (only differing in the random seed for the network initialization); (ii) MLP-k-ens-n is an ensemble of $n$ $k$-layer MLPs, each identically trained but with a unique random network initialization; (iii) MLP-k-no-noise is the same as MLP-k, but is trained without adding any noise on the training data (recall that we train with 3% additive noise by default); and (iv) lin-reg is a linear regression model.
@@ 12 Experimental Results
13.1 Results on Synthetic Data
For our proposed synthetic dataset (cf.
14 sec:synth-data
) we have access to ground truth COTs, and thus in Table 2 we report the RMSE between the predicted outputs and corresponding ground truths on unseen test data. Training on data with artificially added noise improves model robustness significantly. Ensembling only marginally improves performance. Also, MLPs significantly outperform linear regression models.
As mentioned in
15 sec:synth-data
, our synthetic dataset covers five main surface profiles: vegetation, rock, non-photosynthetic vegetation, water and soil. Also recall that vegetation is by far the most common main profile in the dataset ($70.5$% of the data), whereas rock is the second most common one ($10.7$% of the data). To investigate how the prediction accuracy of the best model MLP-5-ens-10 performs on these five different main profiles, we also compute five main profile-conditioned test results (averaged across the different noise percentages, cf. bottom row of Table 2). We found that the RMSEs were 1.45, 3.97, 2.64, 3.44, and 2.42 for vegetation, rock, non-photosynthetic vegetation, water and soil, respectively. Thus the model performs best on vegetation data points, and worst on data points that depict rock.
15.1 Results on Real KappaZeta Data
16 sec:synth-data
@@ 15.1 Results on Real KappaZeta Data
. Such a model is denoted MLP-k-tune. This comparison allows us to quantify how much is gained in prediction accuracy if one trains on the target data domain (the training set of KappaZeta), as opposed to training only on off-domain data (our synthetic dataset). We also compare our MLP models with a U-net segmentation network [^3] We use the open-source FCN8 model [30]. for the same task. This comparison allows us to quantify how much is gained in prediction accuracy if one uses a model that takes into account spatial connectivity among pixels – we expect the U-net to be superior, since U-nets incorporate and take advantage of spatial patterns in images, while the MLPs inherently cannot.
@@
The results are shown in Table 3, where two metrics are reported for each cloud category: F1-score (the harmonic mean of precision and recall) and mean intersection-over-union (mIoU). Both metrics are bounded in the $[0,1]$-range (higher is better). For both metrics we also show the average over the different cloud types. We see that the best MLP-approach (MLP-5-ens-10-tune) obtains a lower mIoU and F1-score compared to the U-net ($0.47$ vs $0.54$ and $0.52$ vs $0.66$, respectively). We re-emphasize however that the MLP approaches perform independent predictions per pixel, whereas the U-net takes into consideration spatial context. Furthermore, we see that model fine-tuning on the KappaZeta training set only slightly improves results on the test set (column 4 vs column 3), which shows that our synthetic dataset can be used to train models which generalize well to real data. It can also be seen that model ensembling only marginally improves results for KappaZeta-tuned models (column 4 vs column 2).
@@ 17.1 Results on Real Data from the Swedish Forest Agency
The Swedish Forest Agency (SFA) is a national authority in charge of forest-related issues in Sweden. Their main function is to promote management of Sweden’s forests that enable the objectives of forest policies to be attained. Among other things, the SFA runs change detection algorithms on satellite imagery, e.g. to detect if forest areas have been felled. For this, they rely on ESA’s scene classification layer (SCL), whose estimates include clouds. The SFA’s analyses require cloud-free images, but the SCL layer is not always accurate enough. Therefore, we applied models that were trained on our synthetic dataset on the SFA’s data in order to classify a set of images as ’cloudy’ or ’clear’.
@@
18 Conclusions
In this work we have introduced a novel synthetic IPA dataset for the task of cloud optical thickness (COT) estimation. Several ML approaches were explored on this data, and it was found that ensembles of MLPs perform best. Despite the fact that our proposed synthetic dataset (and thus associated models) is inherently pixel-independent, the models show promising results on real satellite imagery. In particular, our MLP approach trained on pixel-independent synthetic data can seamlessly transition to real datasets without requiring additional training. To showcase this, we directly applied (without further training) this MLP approach on the task of cloud classification on a novel real image dataset, and achieved an F1-score that is on par with a widely used deep learning model that was explicitly trained on this data. Furthermore, our approach is superior to the ESA-SCL at classifying satellite imagery as clear or cloudy, and can also flexibly generate cloud type segmentation masks via COT thresholding. Code, models and our proposed datasets have be made publicly available at https://github.com/aleksispi/ml-cloud-opt-thick.
Acknowledgments: This work was funded by Vinnova (Swedish Space Data Lab 2.0, grant number 2021-03643), the Swedish National Space Agency and the Swedish Forest Agency.

### v3

@@ 17.2 Fine-tuning with Weaker Labels
Let $\tau^{\text{semi}}$ and $\tau^{\text{opaque}}$ (where $0<\tau^{\text{semi}}<\tau^{\text{opaque}}$) denote the semi-transparent and opaque COT thresholds, respectively. Then, we refine a model using a loss $\mathcal{L}$ which satisfies the following criteria. If $p$ denotes the prediction of a pixel which is labeled as
  • ‘clear’, then
    $\mathcal{L}(p)=\begin{cases}0&\text{ if }p\leq\tau^{\text{semi}},\\ \frac{1}{2}(p-\tau^{\text{semi}})^{2}&\text{ if }p>\tau^{\text{semi}};\end{cases}$
  • ‘opaque cloud’, then
    $\mathcal{L}(p)=\begin{cases}0&\text{ if }p\geq\tau^{\text{opaque}},\\ \frac{1}{2}(p-\tau^{\text{opaque}})^{2}&\text{ if }p<\tau^{\text{opaque}};\end{cases}$
  • ‘semi-transparent cloud’, then
    $\mathcal{L}(p)=\begin{cases}0&\text{ if }\tau^{\text{semi}}\leq p\leq\tau^{\text{opaque}}\\ \frac{1}{2}(p-\tau^{\text{semi}})^{2}&\text{ if }p<\tau^{\text{semi}},\\ \frac{1}{2}(p-\tau^{\text{opaque}})^{2}&\text{ if }p>\tau^{\text{opaque}}.\end{cases}$
19 Experimental Results
20 sec:models
. For COT estimation, we evaluate the following models:
  • MLP-k is a $k$-layer MLP (the main model has $k=5$, cf. Fig. 10), where results are averaged among 10 identical $k$-layer MLP (only differing in the random seed for the network initialization);
  • MLP-k-ens-n is an ensemble of $n$ $k$-layer MLPs, each identically trained but with a unique random network initialization;
  • MLP-k-no-noise is the same as MLP-k, but is trained without adding any noise on the training data (recall that we train with 3% additive noise by default);
  • lin-reg is a linear regression model.
20.1 Results on Synthetic Data
For our proposed synthetic dataset (cf.
21 sec:synth-data
) we have access to ground truth COTs, and thus in Table 2 we report the mean absolute error (MAE) between the predicted outputs and corresponding ground truths on unseen test data. We see that training on data with artificially added noise improves model robustness significantly, whereas ensembling only marginally improves performance. Also, MLPs significantly outperform linear regression models.
Table 2: Mean absolute errors (MAEs) on different variants of the test set of our synthetic dataset. Test-xrefers to the test set with $x$added noise. Ensemble methods marginally improve over single-model ones. Models trained with additive input noise yield significantly better average results. Linear regression performs worst by far. Note that for the single-model variants MLP-5, MLP-5-no-noise and Lin-reg, we show the mean MAE over 10 different network parameter initializations, and standard deviations.
@@ 20.1 Results on Synthetic Data
As mentioned in
22 sec:synth-data
, our synthetic dataset covers five main surface profiles: vegetation, rock, non-photosynthetic vegetation, water and soil. Also recall that vegetation is by far the most common main profile in the dataset ($70.5$% of the data), whereas rock is the second most common one ($10.7$% of the data). To investigate how the prediction accuracy of the best model MLP-5-ens-10 performs on these five different main profiles, we also compute five main profile-conditioned test results (averaged across the different noise percentages, cf. bottom row of Table 2). We found that the MAEs were 1.45, 3.97, 2.64, 3.44, and 2.42 for vegetation, rock, non-photosynthetic vegetation, water and soil, respectively. Thus the model performs best on vegetation data points, and worst on data points that depict rock.
22.1 Results on Real KappaZeta Data
23 sec:synth-data
Table 3: Results on the KappaZeta test set. MLP approaches do not perform as well as U-nets, as is expected since U-nets integrate information from other pixels, which MLPs inherently cannot. See also qualitative results in Fig. 3. Among the MLP approaches, model fine-tuning on the KappaZeta training set only slightly improves results (column 4 vs 3), i.e. our synthetic dataset allows for model generalization to real data, while ensembling tuned models negligibly improves results (column 4 vs 2).
@@ 22.1 Results on Real KappaZeta Data
. Such a model is denoted MLP-k-tune. This comparison allows us to quantify how much is gained in prediction accuracy if one trains on the target data domain (the training set of KappaZeta), as opposed to training only on off-domain data (our synthetic dataset). We also compare our MLP models with a U-net segmentation network [^3] We use the open-source FCN8 model [30]. for the same task. Different to an MLP, a U-net is a fully convolutional architecture that explicitly incorporates contextual information from surrounding pixels in an image when performing predictions. Thus this comparison with a U-net allows us to quantify how much is gained in prediction accuracy if one uses a model that takes into account spatial connectivity among pixels – we expect the U-net to be superior, since U-nets incorporate and take advantage of spatial patterns in images, while the MLPs inherently cannot.
The results are shown in Table 3, where two metrics are reported for each cloud category:
  • F1-score, i.e. the harmonic mean of precision and recall, where precision is the fraction of pixels that were correctly predicted as the given category, and recall is the fraction of pixels of the given category that were predicted as that category;
  • intersection-over-union (IoU), i.e. the ratio between the number of pixels in the intersection and union, respectively, of the predicted and ground truth cloud masks.
Both metrics are bounded in the $[0,1]$-range (higher is better). For both metrics we also show the average result over the different cloud types (these are respectively denoted F1-avg and mIoU), where the average is computed uniformly across the different cloud types, i.e. each category contributes equally to the mean, independently of how common the category is in the dataset.
@@
We see that the best MLP-approach (MLP-5-ens-10-tune) obtains a lower mIoU and F1-score compared to the U-net ($0.47$ vs $0.54$ and $0.52$ vs $0.66$, respectively). We re-emphasize however that the MLP approaches perform independent predictions per pixel, whereas the U-net takes into consideration spatial context. Furthermore, we see that model fine-tuning on the KappaZeta training set only slightly improves results on the test set (column 4 vs column 3), which shows that our synthetic dataset can be used to train models which generalize well to cloud detection on real data. It can also be seen that model ensembling only marginally improves results for KappaZeta-tuned models (column 4 vs column 2).
@@ 24.1 Results on Real Data from the Swedish Forest Agency
The Swedish Forest Agency (SFA) is a national authority in charge of forest-related issues in Sweden. Their main function is to promote management of Sweden’s forests that enable the objectives of forest policies to be attained. Among other things, the SFA runs change detection algorithms on satellite imagery, e.g. to detect if forest areas have been felled. For this, they rely on ESA’s scene classification layer (SCL), which also includes a cloud probability product. The SFA’s analyses require cloud-free images, but the SCL layer is not always accurate enough. Therefore, we applied models that were trained on our synthetic dataset on the SFA’s data in order to classify a set of images as ’cloudy’ or ’clear’.
@@
25 Conclusions
In this work we have introduced a novel synthetic IPA dataset that can be used to train models for predicting cloud types of pixels (e.g. clear, semi-transparent, opaque clouds). In order to enable broad application, in our dataset we chose to use cloud optical thickness (COT) as a proxy for cloud masking. Several ML approaches were explored on this data, and it was found that ensembles of MLPs perform best. Despite the fact that our proposed synthetic dataset (and thus associated models) is inherently pixel-independent, the models show promising results on real satellite imagery. In particular, our MLP approach trained on pixel-independent synthetic data can seamlessly transition to real datasets without requiring additional training. To showcase this, we directly applied (without further training) this MLP approach on the task of cloud classification on a novel real image dataset, and achieved an F1-score that is on par with a widely used deep learning model that was explicitly trained on this data. Furthermore, our approach is superior to the ESA scene classification layer at classifying satellite imagery as clear or cloudy, and can flexibly generate cloud type segmentation masks via COT thresholding. Code, models and our proposed datasets have be made publicly available at https://github.com/aleksispi/ml-cloud-opt-thick.
Acknowledgments: This work was funded by VINNOVA (Swedish Space Data Lab 2.0, grant number 2021-03643, and Swedish Space Data Lab 3.0, grant number 2023-02787). We thank the Swedish Forest Agency for providing the real dataset used in this work.